# Growing Mathlib: maintenance of a large scale mathematical library

### v1

@@ Abstract
Keywords:
Formal mathematics Library development Linting.
@@ 2 Related work
Various interactive theorem provers have come to enjoy large libraries of research mathematics—each with its own strengths and weaknesses. The Mizar Mathematical Library [2] is the oldest and perhaps the largest library (at 3.7 millions lines of code (MLOC)) of formalised mathematics. The Isabelle/HOL system has very powerful automation; the Archive of Formal Proofs (AFP; 4.8 MLOC [^1] The AFP has a broader focus [3], with about half its current articles covering the analysis and verification of particular algorithms. Arguably, this makes the MML a larger library of formalised mathematics.) still has the strongest complex analysis library. The Rocq theorem prover allow expressing complex mathematical objects (such as schemes, the tangent space of a manifold or vector bundles) more naturally, thanks to its foundations including dependent types. While its mathematical components library [22] is significantly smaller (150 000 lines of code), Rocq has been used to formalise major results such as the Four Colour and the Feit–Thompson theorems (which are each of similar size as the mathematical components). Let us also mention the Flyspeck project [15] as an example of combining large libraries in two different proof assistants.
The MML introduced mandatory review of contributions in 2006, as well as tooling to check for duplicate lemmas. (Such tooling also exists for Rocq [10] and Lean [30].) Klein et al have written [4] about maintaining large scale Isabelle libraries. They name many challenges similar to ours, including coordination and collaboration in large and diverse projects. The Isabelle system also employs custom linters [26] — which are not mandated for existing entries, but enforced for new submissions. Zimmermann [38] has written about recent changes to Rocq’s development processes (including switching to github and a pull-request based model, dealing with breaking changes and the challenges with distributing reviews). There is also custom tooling to support Rocq maintainers [39].
@@ 3 The deprecation system
The spectrum of possible breaking changes is rather diverse, and ranges from one-off changes when upgrading to the next Lean version to adjusting to particular refactorings in Mathlib. In principle, Lean’s excellent meta-programming facilities allow writing custom migrations for virtually all changes, transforming working code from one version to the next. Theoretically, such migrations could be arbitrarily complicated. In practice, however, the vast majority of adjustments are very simple to deal with: renaming (the declaration which states) a theorem or definition, renaming files (which requires adjusting sec:linters); these can also help with semantic changes. Lean has a deprecation system specifically designed to handle renamings gracefully. When a declaration is renamed, for a grace period of several months, we add a deprecation marker documenting the rename. This connects the old and new names (and allows providing a custom message to guide the migration process). Upon using a deprecated declaration, the user sees a warning, indicating which declaration to use instead. This allows for a much smoother migration. Dealing with renamed files is harder. Lean looks at file imports before the deprecation system (which is implemented as a linter) has the chance to run—so a file that imported the file with the old name would stop building entirely. The solution is similar: the module with the old name is not erased, but is preserved and stained. Thus, it can be imported without breaking the build—and a linter suggests to replace the old module name, with appropriate imports. These could be either the direct imports of that module, in the case of a reorganization of content; or the renamed module, in the case of a simple move. This informs users of where the declarations that were present in the old file may have ended up.
@@ 4 Semantic linters
and in fact Mathlib uses linters for all of these purposes. To be precise, Mathlib uses two different linter frameworks, with distinct strengths and weaknesses. Lean itself provides a https://lean-lang.org/doc/reference/latest//Notations-and-Macros/Defining-New-Syntaxsyntax of the command as input and can run arbitrary code. [^3] Technically, a linter function returns a , i.e. the only interesting outcomes of the linter are its side effects. Lean’s community-maintained extended library Batteries provides a parallel .Tactic.Linter structure that is similar to the Lean-native linters, but takes a declaration name as its input and runs code in a different monad. For the purpose of the present discussion, the main difference between the two kinds of linters is that the former is designed to run on every command and has access to the syntax tree of each command, while the latter is designed to run once on the “final” environment and only inspects the declarations present in the Environment. We refer to the former linters as “syntax linters” and to the latter linters as “environment linters”. Both are semantic linters, in the sense that they can access the fully elaborated information and not merely the typed syntax. Since both frameworks allow running arbitrary code, the distinction between them is mostly on what they are expected to do and what they can do easily, not on what they could do. Environment linters are good for global validation (e.g. checking lemmas are in normal form), but only warn you after the fact. Syntax linters warn you right away and are best suited for local validation (e.g. the valid notation $\lambda$ x => x is deprecated in favour of fun x $\mapsto$ x). At the time of writing, there are 26 syntax linters defined in Mathlib and 10 more coming from its dependencies, as well as 17 environment linters, all but two defined in Batteries. We mostly focus on the syntax linters: virtually all the environment linters have been described previously [36]. [^4] With the advent of Lean 4, the need for the typeclass linters described there has mostly disappeared: typeclass inference was substantially redesigned in Lean 4 [33], taking into account the insights from Mathlib development. This means some linter warnings have become compiler errors, and the others have become less relevant and did not get reimplemented in Lean 4. To summarize more coherently how Mathlib uses linters, let us group them into general categories and highlight a few examples from each category.
@@ 4.1 Style linters
Style linters primarily inspect the syntax tree of each command. A slightly imprecise but descriptive way of thinking about them is that they match syntax-aware regular expressions. An example in this category is the docstring linter. This performs a simple check on doc-strings, making sure that they begin with exactly one space or line break before the first non-whitespace character, and they perform a similar check for the closing of doc-strings. Uniformizing the appearance of doc-strings contributes to a homogeneous file layout which is helpful when scanning the source code by eye. Another style linter is the itemize contains a valid copyright header, does not import files like Mathlib.lean or Mathlib/Tactic.lean, contains a module doc-string, summarising the content of the file (unless the file is an itemize This ensures a common format for all the Mathlib files and also takes a step in the direction of minimizing the dependency structure of Mathlib. There are other, specific tools that are geared specifically towards controlling imports, such as DirectoryDependency, but this is a quick and simple syntactic check.
@@ 4.1.2 Tracking issues
Certain commands are expected to be rarely used, or are only temporary adaptations. For instance, increasing the time limit for type-checking is something that may be necessary for an individual declaration, but increasing the limit for a whole file is likely a symptom of some deeper issue that should be addressed. Hence, a command such as setOption linter does exactly that: it flags uses of set_option maxHeartbeats x in … instead. Similar linters exist for globalAttributeIn linter that warns that the scope of certain attributes is not “local”, even in the presence of the (usually) scope-limiting Maintenance and code readability The linters in this category are specifically geared to improve maintainability of code and readability. Since Mathlib evolves very quickly, code breakage is fairly common; any tooling which can diagnose and prevent such breakages is immensely helpful. The linters in this section try to be pro-active about code breakage by reducing code brittleness.
@@
The simp, aesop) apply to a wide variety of goals and return similarly varied processed goals; “rigid” tactics (such as apply) perform some very specific operations and may stop working even with very small changes of the current goal state. The simp lemma allows further simplification, making the following ring or Handling multiple side-goals In terms of goal management and readability, the ·. The code style enforced by this linter makes both the number of side-goals that remain after each tactic application as well as which parts of the code deal with which part of the proof completely transparent. This helps with maintenance, since it limits code breakage to easily identified branches, rather than a generic total failure. At the same time, it also improves code readability and structure, since the various arms of the proofs are visually differentiated.
@@
Finally, let us mention the section commands that do not have a matching section, producing a dangling one. As a consequence, the scope of the currently non-Helpful and pedagogical suggestions Some linters attempt to preempt possible misunderstanding or common pitfalls. The have when introducing a non-let when introducing an assumption in have for papercut linter. One common mistake is accidentally using subtraction of natural numbers (which truncates at
@@ 5.2 Speed and coercions
The coercion system is another commonly used component of Lean. Essentially, when a term A, where type t. The coercion is found by synthesizing an instance of the typeclass CoeT should not be declared directly, but rather chained together from the CoeOut, CoeTail typeclasses. The rules for which typeclass to instantiate for a particular coercion are subtle and not universally understood; they are notably more complex compared to Lean 3. As a consequence, a seemingly acceptable instance can cause a large performance regression. For example, B to be a free variable, but doing so causes large performance regressions since such an instance always applies. [^6] https://github.com/leanprover-community/mathlib4/pull/22950 In a similar vein to Subsection 5.1, large hierarchies extending coercion typeclasses should be avoided. Mathlib uses the MonoidHomClass used to extend the FunLike hierarchy. Even worse, since Monoid instances, this also triggers a search through the algebraic hierarchy. The solution was to unbundle inheritance from FunLike and https://github.com/leanprover-community/mathlib4/pull/8386refactor by Anne Baanen performing this unbundling resulted in a 33% speedup for typeclass instance synthesis and an overall 19% decrease in build instructions.
@@ 5.3 Speed and definitional equality
Type checking, and definitional equality in particular, is commonly invoked in the Lean elaboration and kernel checking process and can result in arbitrarily complex computations: terms can be large and their size can increase further when unfolding definitions referenced by the term. Equality checks are intended to be “cheap” basic steps in implementing more complex routines such as typeclass instance synthesis, so slowdowns in checking have a large impact. Generally, definitions are designed to not require unfolding unless the unfolding is requested. Each definition in Lean has a transparency level controlling the circumstances in which it is unfolded. Most tactics are designed to only unfold definitions explicitly marked as reducible, with some tactics also unfolding at the default semireducible transparency level. Marking definitions as irreducible prevents unfolding altogether, skipping a potentially expensive equality check if it does not obviously succeed. A particular example of complex definitional equality checking can be found in structure declarations. Large structures can already be quite complex terms in themselves, which can be hidden behind further definitions and constructions, and the possibility of diamond inheritance can then result in nonobvious equality checks. This all results in high levels of unfolding being required to verify the equality of terms. Typeclasses in Lean are implemented as structures, so the complex typeclass hierarchy also has an impact here. The fast_instance term elaboration macro that normalizes instance declarations by unfolding definitions and replacing subterms with synthesized instances. An initial implementation of this macro resulted in a 5% decrease in type checking time and halved compilation times for many files. [^7] https://github.com/leanprover-community/mathlib4/pull/20993
@@ 5.4 Speed and tactics
During the compilation of Mathlib, Lean spends approximately 14% of its time running the discrimination trees [24, 29] that allow for efficient association of terms with values. A discrimination tree is a tree datastructure whose branches are indexed by symbols (either a constant, or the special symbol *. Still, the approximation is much finer than previously in Lean 3, where terms were indexed by only the first constant that appeared [28]. The simplifier tries to repeatedly rewrite using a set of lemmas and tactics (called a simp from the default large simp only keywords) since many fewer lemmas have to be tested for applicability. Going from simp lowers the maintenance burden, since a change to the default simp only call. The div_self a : a / a = 1 applies if Group, but a field does not form a multiplicative group. Therefore, the https://github.com/leanprover-community/mathlib4/pull/21326 Another avenue to optimizing the use of tactics is to use weaker, faster tactics instead of more powerful but slower tactics. The Aesop tactic provides general proof search [20] and is also invoked as part of other, more specialized tactics to discharge goals. The general nature of this proof search makes Aesop very powerful, but this power is often unnecessary when the proof it needs to find has a predetermined format. We have found that replacing powerful Aesop-based tactics like fun_prop is almost always possible and reduces running time appreciably. Similarly, small tweaks such as applying local hypotheses or trying reflexivity to discharge an equality goal before starting Aesop can lead to appreciable speedups. [^8] https://github.com/leanprover-community/mathlib4/pull/21330
@@ 6 Tracking and addressing technical debt
This concept is all the more important in a large foundational library like Mathlib, especially in light of its cohesive and tighly integrated nature. In this section we describe various types of technical debt we have encountered in Mathlib, and how the community tracks and addresses them.
@@
From 2021–2023, the community ported Mathlib from Lean 3 to Lean 4. This process ought to be documented in more detail elsewhere, but for the purposes of the present discussion we summarize the points of interest. For all practical purposes, Lean 4 is a distinct language from Lean 3—there is no backwards compatibility. Mario Carneiro wrote a tool mathport that automatically translates Lean 3 code to Lean 4 code on a best-effort basis, and a small group of people implemented the common tactics in Lean 4. In October/November 2022, this preparatory work transitioned into a wide-scale and coordinated effort to port the theory files of Mathlib. Over the course of 10 months, the community ported on the order of a million lines of code, or roughly a dozen files per day. Naturally, there were many places where a faithful translation was not possible. For the sake of speed, the porting team decided to liberally annotate issues with “porting notes” that described the problem and the workaround. As a result, by the summer of 2023, the porting process was complete, but Mathlib 4 contained well over 5000 porting notes. Pietro Monticone made a valiant effort to organize and classify the porting notes in February and March of 2024. Subsequently, Anne Baanen, Kim Morrison, and Johan Commelin processed and addressed over two thirds of the porting notes in the first quarter of 2025. Many of these porting notes became obsolete because of new and expanded functionality in Lean 4. At the time of writing, roughly 1500 porting notes remain.
@@
New versions of Lean 4 are released on a monthly basis. Mathlib maintains a branch that tracks the nightly releases, which simplifies migrating to the new version of Lean 4 when it is released. This tracking branch provides feedback for the development of Lean 4: if the daily transition to a new nightly version of Lean 4 requires some unsatisfactory workarounds, then these are annotated with #adaptation_note comments. In essence, these notes have the same functionality as porting notes. Occasionnaly, there is no simple solution for a problem that arises from a new version of Lean 4. In such an event the old feature can be re-enabled via a backwards compatibility flag. This allows the Mathlib community to transition to the new version of Lean 4 at a reasonable pace.
@@
The shell script scripts/technical-debt-metrics.sh is used to create a report of the current state of known technical debt in Mathlib. This includes tracking adaption notes, porting notes and backwards compatibility flags. Continuous integration runs this script on a weekly basis, and posts the report (including the delta with respect to the previous week) to the Zulip chat of the Lean community. Additionally, this report is computed for each pull request, and posted to a dedicated “PR summary” comment at the top of the pull request. This allows the reviewer to quickly see the current state of technical debt in the pull request, and take this into account when deciding whether to approve the pull request. Mathlib also has a mechanism in place to surface pull requests addressing technical debt: this includes both a dedicated label and triage dashboard (see Section 7.2) and a Zulip stream where such PRs can be brought to attention. One downside of tracking and reporting technical debt in this manner is indicated by Goodhart’s Law: “When a measure becomes a target, it ceases to be a good measure.” In this context, this means that the community is inclined to focus on the technical debt that is being tracked, whereas there may be other forms of technical debt that fly under the radar. To alleviate this, there is a dedicated thread on the Lean community’s Zulip chat listing less measurable forms of technical and organisational debt. (One example of the latter is documenting unwritten informal Mathlib policy: sharing such knowledge enables other contributors to learn this for themselves and lowers the barrier towards contribution. An example of the former is reorganizing top-level directories: due to Mathlib’s age, some of these are not coherent any more.) Everybody can add items to the list, facilitating a collective understanding of future maintenance tasks.

### v2

@@ 2 Related work
Various interactive theorem provers have come to enjoy large libraries of research mathematics—each with its own strengths and weaknesses. The Mizar Mathematical Library [2] is the oldest and perhaps the largest library (at 3.7 millions lines of code (MLOC)) of formalised mathematics. The Isabelle/HOL system has very powerful automation; the Archive of Formal Proofs (AFP; 4.8 MLOC [^1] The AFP has a broader focus [3], with about half its current articles covering the analysis and verification of particular algorithms. Arguably, this makes the MML a larger library of formalised mathematics.) still has the strongest complex analysis library. The Rocq theorem prover allows expressing complex mathematical objects (such as schemes, the tangent space of a manifold or vector bundles) more naturally, thanks to its foundations including dependent types. While its mathematical components library [23] is significantly smaller (150 000 lines of code), Rocq has been used to formalise major results such as the Four Colour and the Feit–Thompson theorems (which are each of similar size as the mathematical components). Let us also mention the Flyspeck project [15] as an example of combining large libraries in two different proof assistants.
The MML introduced mandatory review of contributions in 2006, as well as tooling to check for duplicate lemmas. (Such tooling also exists for Rocq [10] and Lean [31].) Klein et al have written [4] about maintaining large-scale Isabelle libraries. They name many challenges similar to ours, including coordination and collaboration in large and diverse projects. The Isabelle system also employs custom linters [27] — which are not mandated for existing entries, but enforced for new submissions. Zimmermann [39] has written about recent changes to Rocq’s development processes (including switching to GitHub and a pull-request based model, dealing with breaking changes and the challenges with distributing reviews). There is also custom tooling to support Rocq maintainers [40]. Huch’s in-progress PhD thesis [17] describes efforts to address scaling issues in the AFP, including a discussion of maintainenance efforts.
@@ 3 The deprecation system
The spectrum of possible breaking changes is rather diverse, and ranges from one-off changes when upgrading to the next Lean version to adjusting to particular refactorings in Mathlib. In principle, Lean’s excellent meta-programming facilities allow writing custom migrations for virtually all changes, transforming working code from one version to the next. Theoretically, such migrations could be arbitrarily complicated. In practice, however, the vast majority of adjustments are very simple to deal with: renaming (the declaration which states) a theorem or definition, renaming files (which requires adjusting sec:linters); these can also help with semantic changes.
Lean has a deprecation system specifically designed to handle renamings gracefully. When a declaration is renamed, for a grace period of several months, we add a deprecation marker documenting the rename. This connects the old and new names (and allows providing a custom message to guide the migration process). Upon using a deprecated declaration, the user sees a warning, indicating which declaration to use instead. This allows for a much smoother migration.
Dealing with renamed files is harder. Lean looks at file imports before the deprecation system (which is implemented as a linter) has the chance to run—so a file that imported the file with the old name would stop building entirely. The solution is similar: the module with the old name is not erased, but is preserved and stained. Thus, it can be imported without breaking the build—and a linter suggests to replace the old module name, with appropriate imports. These could be either the direct imports of that module, in the case of a reorganization of content; or the renamed module, in the case of a simple move. This informs users of where the declarations that were present in the old file may have ended up.
@@ 4 Semantic linters
and in fact Mathlib uses linters for all of these purposes.
To be precise, Mathlib uses two different linter frameworks, with distinct strengths and weaknesses. Lean itself provides a https://lean-lang.org/doc/reference/latest//Notations-and-Macros/Defining-New-Syntaxsyntax of the command as input and can run arbitrary code. [^3] Technically, a linter function returns a , i.e. the only interesting outcomes of the linter are its side effects.Lean’s community-maintained extended library Batteries provides the parallel structure
command
and has access to the syntax tree of each command, while the latter is designed to run once on the “final” environment and only inspects the declarations present in the Environment. We refer to the former linters as “syntax linters” and to the latter linters as “environment linters”. Both are semantic linters, in the sense that they can access the fully elaborated information and not merely the typed syntax. Since both frameworks allow running arbitrary code, the distinction between them is mostly on what they are expected to do and what they can do easily, not on what they could do. Environment linters are good for global validation (e.g. checking lemmas are in normal form), but only warn you after the fact. Syntax linters warn you right away and are best suited for local validation (e.g. the valid notation $\lambda$ x => x is deprecated in favour of fun x $\mapsto$ x).
At the time of writing, there are 26 syntax linters defined in Mathlib and 10 more coming from its dependencies, as well as 17 environment linters, all but two defined in Batteries. We mostly focus on the syntax linters: virtually all the environment linters have been described previously [37]. [^4] With the advent of Lean 4, the need for the typeclass linters described there has mostly disappeared: typeclass inference was substantially redesigned in Lean 4 [34], taking into account the insights from Mathlib development. This means some linter warnings have become compiler errors, and the others have become less relevant and did not get reimplemented in Lean 4. To summarize more coherently how Mathlib uses linters, let us group them into general categories and highlight a few examples from each category.
@@ 4.1 Style linters
Style linters primarily inspect the syntax tree of each command. A slightly imprecise but descriptive way of thinking about them is that they match syntax-aware regular expressions.
An example in this category is the docstring linter. This performs a simple check on doc-strings, making sure that they begin with exactly one space or line break before the first non-whitespace character, and they perform a similar check for the closing of doc-strings.
Uniformizing the appearance of doc-strings contributes to a homogeneous file layout which is helpful when scanning the source code by eye.
Another style linter is the itemize
contains a valid copyright header,
does not import files like Mathlib.lean or Mathlib/Tactic.lean,
contains a module doc-string, summarising the content of the file (unless the file is an itemize This ensures a common format for all the Mathlib files and also takes a step in the direction of minimizing the dependency structure of Mathlib. There are other, specific tools that are geared specifically towards controlling imports, such as DirectoryDependency, but this is a quick and simple syntactic check.
@@ 4.1.2 Tracking issues
Certain commands are expected to be rarely used, or are only temporary adaptations. For instance, increasing the time limit for type-checking is something that may be necessary for an individual declaration, but increasing the limit for a whole file is likely a symptom of some deeper issue that should be addressed. Hence, a command such as setOption linter does exactly that: it flags uses of set_option maxHeartbeats x in … instead.
Similar linters exist for globalAttributeIn linter that warns that the scope of certain attributes is not “local”, even in the presence of the (usually) scope-limiting Maintenance and code readability
The linters in this category are specifically geared to improve maintainability of code and readability. Since Mathlib evolves very quickly, code breakage is fairly common; any tooling which can diagnose and prevent such breakages is immensely helpful. The linters in this section try to be pro-active about code breakage by reducing code brittleness.
@@
The simp, aesop) apply to a wide variety of goals and return similarly varied processed goals; “rigid” tactics (such as apply) perform some very specific operations and may stop working even with very small changes of the current goal state.
The simp lemma allows further simplification, making the following ring or Handling multiple side-goals In terms of goal management and readability, the ·. The code style enforced by this linter makes both the number of side-goals that remain after each tactic application as well as which parts of the code deal with which part of the proof completely transparent. This helps with maintenance, since it limits code breakage to easily identified branches, rather than a generic total failure. At the same time, it also improves code readability and structure, since the various arms of the proofs are visually differentiated.
@@
Finally, let us mention the section commands that do not have a matching section, producing a dangling one. As a consequence, the scope of the currently non-Helpful and pedagogical suggestions
Some linters attempt to preempt possible misunderstanding or common pitfalls. The have when introducing a non-let when introducing an assumption in have for papercut linter. One common mistake is accidentally using subtraction of natural numbers (which truncates at
@@ 5.1 Speed and typeclasses
Mathlib relies on the Lean typeclass system to perform automatic inference of many implicit arguments. Typeclasses are flexible and have low infrastructure requirements on user code, but the flexibility also means designs have to be carefully consider to optimize their tradeoffs. As the typeclass hierarchy grows, synthesis of instances becomes slower since a larger search space needs to be explored. It is not only important to ensure synthesis succeeds quickly: if synthesis fails, tactics can decide to fall back on an alternative proof strategy. Lean’s instance synthesis algorithm fails only after exploring the full search space, and it is therefore especially important that the space does not grow too large.
@@
The speed tradeoffs changed when upgrading from Lean 3 to Lean 4. A consequence of the bundled approach to typeclass design is that the number of classes becomes multiplicative when extending along two distinct axes. Taking the LinearOrderedCommMonoid, or in the algebraic hierarchy to give LinearOrderedCommGroup. When looking for a [CommMonoid M] parameter with [PartialOrder M] with https://github.com/leanprover-community/mathlib4/pull/20676refactor by Yuyang Zhao resulted in a 20% decrease in typeclass inference time, resulting in a 6% overall speedup to Mathlib compilation.
@@ 5.2 Speed and coercions
The coercion system is another commonly used component of Lean. Essentially, when a term A, where type t. The coercion is found by synthesizing an instance of the typeclass CoeT should not be declared directly, but rather chained together from the CoeOut, CoeTail typeclasses. The rules for which typeclass to instantiate for a particular coercion are subtle and not universally understood; they are notably more complex compared to Lean 3. As a consequence, a seemingly acceptable instance can cause a large performance regression. For example, B to be a free variable, but doing so causes large performance regressions since such an instance always applies. [^6] https://github.com/leanprover-community/mathlib4/pull/22950
In a similar vein to Subsection 5.1, large hierarchies extending coercion typeclasses should be avoided. Mathlib uses the MonoidHomClass used to extend the FunLike hierarchy. Even worse, since Monoid instances, this also triggers a search through the algebraic hierarchy. The solution was to unbundle inheritance from FunLike and https://github.com/leanprover-community/mathlib4/pull/8386refactor by Anne Baanen performing this unbundling resulted in a 33% speedup for typeclass instance synthesis and an overall 19% decrease in build instructions.
@@ 5.3 Speed and definitional equality
Type checking, and definitional equality in particular, is commonly invoked in the Lean elaboration and kernel checking process and can result in arbitrarily complex computations: terms can be large and their size can increase further when unfolding definitions referenced by the term. Equality checks are intended to be “cheap” basic steps in implementing more complex routines such as typeclass instance synthesis, so slowdowns in checking have a large impact.
Generally, definitions are designed to not require unfolding unless the unfolding is requested. Each definition in Lean has a transparency level controlling the circumstances in which it is unfolded. Most tactics are designed to only unfold definitions explicitly marked as reducible, with some tactics also unfolding at the default semireducible transparency level. Marking definitions as irreducible prevents unfolding altogether, skipping a potentially expensive equality check if it does not obviously succeed.
A particular example of complex definitional equality checking can be found in structure declarations. Large structures can already be quite complex terms in themselves, which can be hidden behind further definitions and constructions, and the possibility of diamond inheritance can then result in nonobvious equality checks. This all results in high levels of unfolding being required to verify the equality of terms.
Typeclasses in Lean are implemented as structures, so the complex typeclass hierarchy also has an impact here. The fast_instance term elaboration macro that normalizes instance declarations by unfolding definitions and replacing subterms with synthesized instances. An initial implementation of this macro resulted in a 5% decrease in type checking time and halved compilation times for many files. [^7] https://github.com/leanprover-community/mathlib4/pull/20993
@@ 5.4 Speed and tactics
During the compilation of Mathlib, Lean spends approximately 14% of its time running the discrimination trees [25, 30] that allow for efficient association of terms with values. A discrimination tree is a tree datastructure whose branches are indexed by symbols (either a constant, or the special symbol *. Still, the approximation is much finer than previously in Lean 3, where terms were indexed by only the first constant that appeared [29].
The simplifier tries to repeatedly rewrite using a set of lemmas and tactics (called a simp from the default large simp only keywords) since many fewer lemmas have to be tested for applicability. Going from simp lowers the maintenance burden, since a change to the default simp only call.
The div_self a : a / a = 1 applies if Group, but a field does not form a multiplicative group. Therefore, the https://github.com/leanprover-community/mathlib4/pull/21326
Another avenue to optimizing the use of tactics is to use weaker, faster tactics instead of more powerful but slower tactics. The Aesop tactic provides general proof search [21] and is also invoked as part of other, more specialized tactics to discharge goals. The general nature of this proof search makes Aesop very powerful, but this power is often unnecessary when the proof it needs to find has a predetermined format. We have found that replacing powerful Aesop-based tactics like fun_prop is almost always possible and reduces running time appreciably. Similarly, small tweaks such as applying local hypotheses or trying reflexivity to discharge an equality goal before starting Aesop can lead to appreciable speedups. [^8] https://github.com/leanprover-community/mathlib4/pull/21330
@@ 6 Tracking and addressing technical debt
This concept is all the more important in a large foundational library like Mathlib, especially in light of its cohesive and tightly integrated nature. In this section we describe various types of technical debt we have encountered in Mathlib, and how the community tracks and addresses them.
@@
From 2021–2023, the community ported Mathlib from Lean 3 to Lean 4. This process ought to be documented in more detail elsewhere, but for the purposes of the present discussion we summarize the points of interest.
For all practical purposes, Lean 4 is a distinct language from Lean 3—there is no backwards compatibility. Mario Carneiro wrote a tool mathport that automatically translates Lean 3 code to Lean 4 code on a best-effort basis, and a small group of people implemented the common tactics in Lean 4. In October/November 2022, this preparatory work transitioned into a wide-scale and coordinated effort to port the theory files of Mathlib. Over the course of 10 months, the community ported on the order of a million lines of code, or roughly a dozen files per day.
Naturally, there were many places where a faithful translation was not possible. For the sake of speed, the porting team decided to liberally annotate issues with “porting notes” that described the problem and the workaround. As a result, by the summer of 2023, the porting process was complete, but Mathlib 4 contained well over 5000 porting notes.
Pietro Monticone made a valiant effort to organize and classify the porting notes in February and March of 2024. Subsequently, Anne Baanen, Kim Morrison, and Johan Commelin processed and addressed over two thirds of the porting notes in the first quarter of 2025. Many of these porting notes became obsolete because of new and expanded functionality in Lean 4. At the time of writing, roughly 1500 porting notes remain.
@@
New versions of Lean 4 are released on a monthly basis. Mathlib maintains a branch that tracks the nightly releases, which simplifies migrating to the new version of Lean 4 when it is released. This tracking branch provides feedback for the development of Lean 4: if the daily transition to a new nightly version of Lean 4 requires some unsatisfactory workarounds, then these are annotated with #adaptation_note comments. In essence, these notes have the same functionality as porting notes. Occasionnally, there is no simple solution for a problem that arises from a new version of Lean 4. In such an event the old feature can be re-enabled via a backwards compatibility flag. This allows the Mathlib community to transition to the new version of Lean 4 at a reasonable pace.
@@
The shell script scripts/technical-debt-metrics.sh is used to create a report of the current state of known technical debt in Mathlib. This includes tracking adaptation notes, porting notes and backwards compatibility flags. Continuous integration runs this script on a weekly basis, and posts the report (including the delta with respect to the previous week) to the Zulip chat of the Lean community. Additionally, this report is computed for each pull request, and posted to a dedicated “PR summary” comment at the top of the pull request. This allows the reviewer to quickly see the current state of technical debt in the pull request, and take this into account when deciding whether to approve the pull request.
Mathlib also has a mechanism in place to surface pull requests addressing technical debt: this includes both a dedicated label and triage dashboard (see Section 7.2) and a Zulip stream where such PRs can be brought to attention.
One downside of tracking and reporting technical debt in this manner is indicated by Goodhart’s Law: “When a measure becomes a target, it ceases to be a good measure.” In this context, this means that the community is inclined to focus on the technical debt that is being tracked, whereas there may be other forms of technical debt that fly under the radar. To alleviate this, there is a dedicated thread on the Lean community’s Zulip chat listing less measurable forms of technical and organisational debt. (One example of the latter is documenting unwritten informal Mathlib policy: sharing such knowledge enables other contributors to learn this for themselves and lowers the barrier towards contribution. An example of the former is reorganizing top-level directories: due to Mathlib’s age, some of these are not coherent any more.) Everybody can add items to the list, facilitating a collective understanding of future maintenance tasks.